OpenStudy (anonymous):

If you have: 3^(x)*5^(x-2) - 3^(4x) = 0 Is this a valid factorization: 3^x(5^(x-2) - 3^(4)) = 0 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

because when you multiply you add the exponents, not multiply them

OpenStudy (anonymous):

\[3^x\times 5^{x-2}-3^{4x}\] you can factor out a \[3^x\] but you get \[3^x(5^{x-2}-3^{3x})\]

OpenStudy (anonymous):

since \[3^x\times 3^{3x}=3^{x+3x}=3^{4x}\]